# Not All Attention is Needed: Parameter and Computation Efficient Tuning for Multi-modal Large Language Models via Effective Attention Skipping

### v2

@@ Abstract
Recently, Multi-modal Large Language Models (MLLMs) have garnered an influx of interest from both academia and industry. However, for the downstream task applications, MLLMs not only require to update a large number of parameters but also consume excessive computation. In this paper, we propose a novel parameter and computation efficient tuning method for MLLMs, termed Effective Attention Skipping (EAS). Concretely, we first reveal that multi-head attentions (MHAs) in MLLMs, the primary source of computation, are often redundant to downstream tasks. Based on this observation, EAS evaluates attention redundancy and skips the less important MHAs to speed up inference. Besides, we also propose a novel propagation-of-information adapter (PIA) to serve the attention skipping while maintaining parameter efficiency. More importantly, PIA can be further re-parameterized into feed-forward networks (FFNs) for zero-extra latency. To validate EAS, we apply it to a recently proposed MLLM called LaVIN, and conduct extensive experiments on a vision-language benchmark, namely ScienceQA. The experimental results show that EAS can not only retain the high performance of LaVIN but also reduce the updated parameters scale greatly while speeding up the inference speed to a large extent. For instance, LaVIN-EAS can obtain 89.98% accuracy while accelerating the inference speed by 2.2 times. Our code is given in https://github.com/DoubtedSteam/EAS
@@ 1 Introduction
Recently, the great success of Large Language Models (LLMs) also sparks an influx of interest in extending these giant models to more modalities, i.e., Multi-modal Large Language Models (MLLMs) [38, 32, 25]. However, compared with the unimodal LLMs, MLLMs inevitably require much more computation with the introduction of new modalities [3, 1, 25, 32], and their adaptions to downstream tasks also take more training and parameter expenditure [22, 28, 27]. In this case, numerous efforts have been recently devoted to the efficient tuning of MLLMs [54, 12, 53].
Among these advancements, Wu et al.[53] propose a non-trivial task for large-scaled pre-trained models, termed parameter and computation efficient transfer learning (PCETL). A key intuition behind PCETL is that the sheer size of model parameters and computation is critical for large-scale pre-training but redundant to specific tasks. In this case, for downstream task adaptions, PCETL not only requires to reduce the number of updated parameters, akin to PETL [16, 17, 47], but also needs to remove the redundant modules for better efficiency. Wu et al.[53] also propose a strong baseline called dynamic architecture skipping (DAS) to evaluate and skip the redundant Transformer layers with lightweight adapters, saving both parameter and computation overhead on downstream tasks.
Figure 1: (a) Running time of LaVIN [38], DAS [53] and our EAS. (b) Performance and speed comparisons of skipping different numbers of MHA and FFN by our EAS on ScienceQA.
@@
Despite the effectiveness, we argue that this strong baseline still leaves ample room to improve. Above all, DAS [53] only considers the entire Transformer layer for redundancy evaluation, and a granular scheme deserves more explorations. As shown in Fig. 2-(a), multi-head attention (MHA) consumes 35% more time than feed-forward network (FFN), while its parameter size is only half. From existing literature [40, 49], we also notice that MHA is mainly in charge of dependency modeling, and FFN is often used to improve model capacity. As shown in Fig. 2-(b), skipping a certain number of MHAs does not affect performance, while it will results in a drastic drop when skipping FFNs. In this case, we think that FFN stores more pre-training knowledge for downstream task adaption. In contrast, for a simple task, the repeated attention modeling may be of limited significance. Based on these observations, we make an argument in this paper:
Not all attention is needed for the parameter and computation efficient tuning of MLLMs.”
Another shortcoming of DAS is its skip connections with adapters, which still incurs non-negligible latency during inference. As revealed in recent progresses [38, 12], the additional branches added by adapters inevitably slow down inference although they only brings some FLOPs[38]. The same as shown in Fig.2, the latency caused by adapters takes up to 37.9% inference time of LaVIN [38]. This drawback somewhat undermines DAS’s efficiency for MLLMs.
@@
To overcome these limitations, we propose a novel approach for the parameter and computation efficient tuning of MLLMs in this paper, termed Effective Attention Skipping (EAS). The principle of EAS is to only skip the redundant attention modeling for better efficiency and performance. To avoid the latency caused by adapters, we also equip EAS with an innovative parameter efficient module called propagation-of-information adapter (PIA), which can not only replace the skipped MHAs for adaption, but also can be seamlessly re-parameterized into FFNs for extra zero cost during inference. With these innovative designs, EAS can help MLLMs achieve better PCETL results on vision-language tasks, i.e., true model acceleration with a small number of trainable parameters.
@@
To validate EAS, we first apply it to a recent MLLM called LaVIN [38], and conduct extensive experiments on the open-set multi-modal question answer benchmark, namely ScienceQA [36]. To align DAS [53], we also apply EAS to a representative VL pre-trained model called METER [10] on three VL benchmarks, i.e., VQA2.0 [13], NLVR2 [45] and Flickr30K [43]. The experimental results show that EAS can not only retain high performance and parameter efficiency against existing PETL and PCETL methods [54, 38, 53], but also significantly speed up inference speed on downstream tasks. For instance, EAS can improve the inference speed by 1.90×1.90$\times$1.90 × to DAS [53] and 2.18×2.18$\times$2.18 × to the default LaVIN [38] without performance degradations.
Overall, our contributions can be summarized as follows:
@@
  • We apply our EAS to a recent MLLM and a representative VLP model, i.e., LaVIN and METER, on four VL benchmarks. The experiments show the obvious merits of EAS in both parameter and computation efficiencies, e.g., speeding up LaVIN by 2.18$\times$ without performance degradations.
@@ 2.1 Vision-language pre-trained models
Recent years have witnessed the rapid development of vision-language pre-trained (VLP) models [29, 35, 44, 6, 51] on various vision-language tasks [14, 46, 36]. To facilitate vision-language alignment and learn generalized multi-modal representations, VLP models also adopt self-supervised objects and conduct pre-training on massive image-text pairs, such as masked language modeling [21, 34], masked image modeling [2, 5, 8] and image-text matching [10, 23].
With the advent of large language models (LLMs) [55, 50, 42], recent advance [38, 32, 25] focus on extending these LLMs to vision-language tasks, i.e., multi-modal large language models (MLLMs). For instance, BLIP-2[26] introduces a QFormer, to bridge the gap between vision and language modalities, selecting relevant visual semantics as additional tokens for LLMs. MINI-GPT4[57] uses a projection layer to map visual features into the language model. LLaVA [33] also follows this principle, but proposes a carefully designed training scheme in addition. Meanwhile, some efforts are also devoted to parameter-efficiently transfer learning of LLMs to vision-language tasks [14, 36]. These methods insert lightweight modules like adapter [54] for downstream adaption instead of fully tuning LLMs [17]. Meanwhile, the use of visual tokens greatly exacerbates the already high computation of MLLMs [32, 54]. In this paper, we focus on both parameter and computation efficient tuning for MLLMs.
@@ 2.2 Parameter and Computation Efficient Tuning
Parameter-efficient transfer learning (PETL) is proposed to save the training and storage costs of LLMs [4, 30, 17]. The principle of PETL is to transfer the pre-trained models to downstream tasks with only a small number of trainable parameters [4, 30], and its methodologies can be roughly divided into three main categories, i.e. prompt tuning [18, 56], adapter [20, 11] and the re-parameterized methods [17, 37]. Very recently, Wu et al.[53] propose a new learning task called parameter and computation efficient transfer learning (PCETL), which further requires to reduce model redundancy based on the setting of PETL. Wu et al.[53] also propose a strong baseline called Dynamic Architecture Skipping (DAS) for PCETL, which evaluates the redundancy of Transformer layers by a reinforcement learning method and replaces the most redundant layers with adapters [47]. Motivated by [17, 53], we also focus on PCETL for MLLMs in this paper. We propose a granular exploration for MLLMs, and address the latency caused by using adapters for skip connections.
Figure 3: Illustrations of the main components of the proposed Effective Attention Skipping (EAS). (a) The architecture of propagation-of-information adapter (PIA). PIA uses a multi-path design for up- and down-samplings, which can help to perform information exchange, like MHA, and modality routing [38] for MLLMs. (b) The deployment of PIA. PIA can serve to replace the skipped MHA as a parameter efficient method for task adaption. After training, its parameters can be re-parameterized into FFN, incurring no extra latency. (c) The process of attention redundancy evaluation. Similar with DAS [53], EAS also adopts a $k$-armed bandit based algorithm for the automatic redundancy evaluation on MHAs of MLLMs. After evaluation, we skip the redundant MHAs with PIAs.
@@ 3 Preliminary
We first recap the principle of parameter and computation efficient transfer learning (PCETL) [53] for multi-modal large language models (MLLMs). Concretely, given an MLLM $G(\cdot|\theta)$, which consists $n$ Transformer layers, and its whole parameters denoted as $\mathbf{\theta}=\{\theta_{1},\theta_{2},...,\theta_{n}\}$, the objective of PCETL can be defined by
@@
From Eq.1, we can see that PCETL needs to reduce parameter cost during adaption, just like PETL [30, 17, 47]. Meanwhile, it also requires practitioners to speed up inference by evaluating and removing the redundant modules. Compared with previous model compression and acceleration tasks [24, 41, 7], one difference of PCETL is that the original model weights should keep unchanged, so as to facilitate more task adaptions, which is essential for the giant models that are expensively pre-trained with massive data.
Wu et al.[53] also propose a strong baseline called dynamic architecture skipping (DAS), which considers the entire Transformer layer for redundancy evaluation. As discussed above, the components in MLLMs often serve different roles, i.e., MHA and FFN. In this case, more detailed evaluations are beneficial for PCETL. Besides, the skip connections in DAS are aided by adapters [16, 47] which also incur non-negligible latency during the inference [53, 17].
@@ 4.1 Overview
Concretely, given an MLLMs $G(\cdot|\theta)$, EAS aims to remove its redundant components for downstream task adaption. Different to DAS [53], the evaluation target of EAS is more specific, i.e., MHAs of a MLLM. Thus, the objective can be defined by
@@
where $\bm{\theta}^{A}_{\mathcal{K}}$ represent a subset of MHA modules and $\bm{\theta}^{F}$ denote all FFN modules. In practice, similar to DAS [53], we first conduct a reinforcement learning based redundancy evaluation on MHAs for downstream tasks. Afterwards, given the redundancy scores, we determine which MHAs to skip, and replace them with an adapter-based connection.
However, adapters will account for excessive inference time, as shown in Fig. 2-(a). The widely used re-parameterized methods like LoRA [17] can approximate the QKV projections but cannot replace the entire MHA for skip connections.
@@ 4.2 Propagation-of-Information Adapter
When the redundant MHAs are identified, the next key step is how to effectively skip them without additional latency. To approach this target, we propose a novel re-paremeterizable adapter for EAS, termed Propagation-of-Information Adapter (PIA).
As shown in Fig. 3-(a), PIA also adopts a bottleneck structure to scale down the hidden dimension of the inputs, akin to existing adapters [47, 37], thereby achieving a low-rank approximation of full tuning [17, 37, 52]. However, it also differs in its inner path designs.
@@
where $f_{d1}(\cdot)$ and $f_{d2}(\cdot)$ are the two linear projections, and $avg(\cdot)$ refers to average pooling. $\mathbf{H}^{(i)}\in\mathbb{R}^{n\times r}$ has a much smaller dimension than $\mathbf{X}^{(i)}$, i.e., $r\ll d$.
@@
During up-sampling, we also use two separate linear projections for the hidden features, and adopt a path router for their weighted combination. Thus, the output features $\mathbf{X}^{(i)^{\prime}}$ is obtained by
@@
$\mathrm{where}\ \ \bm{\alpha}$ $=[\alpha_{1},\alpha_{2}],\ \ \bm{\alpha}=\mathrm{router}(avg(\mathbf{X}^{(i)})).$
Here, $f_{u1}(\cdot)$ and $f_{u2}(\cdot)$ are two linear projections, and $\mathrm{router}(\cdot)$ is the routing function defined by
@@
Overall, the multi-path design can still hold parameter efficiency via low-dimensional projections, acting the role of MHA in information exchange.
Re-parameterization. With the effective structure, PIA also needs to re-parameterize its weights into the models to further improve computation efficiency. Here, we embed PIA into FFN, and avoid the skip connections [53] to retain network complexity.
@@
To re-parameterize PIA into the nearby projection weights of FFN [17, 39], we still need to convert PIA into one linear layer:
@@
Here, $\mathbf{W}_{p}$ is the approximated low-rank weight matrix, and it then can be re-parameterized into the nearby weight matrices of FFN, like [17, 37].
Note that, the average feature in Eq. 4, i.e. $\hat{\mathbf{x}}^{(i)}=avg(\mathbf{X}^{(i)})$, and the path router are example-dependent. In this case, the re-parameterization of PIA will be executed after the first decoding step of MLLMs. Considering the fact that most MLLMs often needs to decode a long sequence [38, 54], PIA can still save massive computation during inference.

### v3

@@ Abstract
Recently, Multi-modal Large Language Models (MLLMs) have garnered an influx of interest from both academia and industry. Despite great progress, MLLMs still have a high demand of downstream task tuning for applications, which consumes excessive parameter and computation overhead. In this paper, we propose a novel parameter and computation efficient tuning method for MLLMs, termed Effective Attention Skipping (EAS). Concretely, we first reveal that multi-head attentions (MHAs) in MLLMs, the primary source of computation, are often redundant to downstream tasks. Based on this observation, EAS evaluates attention redundancy and skips the less important MHAs to speed up inference. Besides, we also propose a novel propagation-of-information adapter (PIA) to serve attention skipping while maintaining parameter efficiency. More importantly, PIA can be further re-parameterized into feed-forward networks (FFNs) for zero-extra latency. To validate EAS, we apply it to three common MLLMs, and conduct extensive experiments on six downstream tasks. The experimental results show that EAS can not only retain the high performance of MLLMs but also reduce the updated parameters scale greatly, while speeding up inference speed to a large extent. For instance, LLaVA-EAS can obtain 96.2% accuracy while accelerating the inference speed by about +20% on ScienceQA. Our code is publicly released at https://github.com/DoubtedSteam/EAS.
keywords:
multi-modal large language models, parameter and computation efficient transfer learning
@@ 1 Introduction
Recently, the great success of Large Language Models (LLMs) also sparks an influx of interest in extending these giant models to more modalities, i.e., Multi-modal Large Language Models (MLLMs) [luo2023cheap, liu2023visual, li2023blip, llava-1.5, Internvl]. However, compared with the unimodal LLMs, MLLMs inevitably require much more computation with the introduction of new modalities [bao2022vlmo, alayrac2022flamingo, li2023blip, liu2023visual]. For instance, compared with LLaMA [touvron2023llama], LLaVA [llava-1.5] takes about 6.2 times the computation overhead to solve the questions from the ScienceQA [lu2022learn]. Moreover, the scale of available data for MLLMs is much smaller than LLMs [zhang2023llama, gao2023llama, wu2023parameter], especially the manually labeled SFT examples. In this case, the generalization of MLLMs is often limited, and the requirement of downstream task adaption is often demanded.
Recently, numerous efforts are devoted to the efficient tuning of MLLMs [conf/iclr/HeZMBN22, conf/cvpr/Sung0B22, conf/iclr/HuSWALWWC22]. Among these advancements, [wu2023parameter] propose a non-trivial task for large-scaled pre-trained models, termed parameter and computation efficient transfer learning (PCETL). A key intuition behind PCETL is that the sheer size of MLLMs’ parameters and computation is critical for large-scale pre-training but redundant to specific tasks. In this case, for downstream task adaptions, PCETL not only requires to reduce the number of updated parameters, akin to PETL [houlsby2019parameter, hu2022lora, sung2022vl, journals/ijcv/ZhaoWZ00S24], but also needs to remove the redundant modules for better efficiency.
Figure 1: (a) Latency comparison between EAS and other PETL methods, including Adapter [sung2022vl] and DAS [wu2023parameter], and the default LLaVA [conf/nips/LiuLWL23a]. (b) Performance and speed comparisons of skipping different numbers of MHA and FFN by our EAS on ScienceQA.
@@
However, achieving the target of PCETL still remains challenging. On the one hand, the redundancy of MLLMs’ components for downstream tasks still requires more in-depth exploration, e.g., multi-head attention (MHA) and feed-forward network (FFN). As noted in existing literature [martins2020sparse, tolstikhin2021mlp], MHA mainly acts a role of dependency modeling, while FFN is designed to enhance model capacity. Furthermore, MHA consumes 35% more time than FFN despite having only half the number of parameters, as shown in Fig. 1-(a). In [wu2023parameter], Wu et al. take the entire transformer layers as a candidate to evaluate redundancy. In this coarse-grained way, only a few candidates can be skipped, limiting the merits of PCETL. Meanwhile, we also observe that skipping a certain number of MHAs does not affect performance, while it will result in a drastic drop when skipping FFNs, as shown in Fig. 1-(b).
On the other hand, how to acheive true speed-up of MLLMs remains an open problem for PCETL. Specifically, although the lightweight adapters used to replace the original module only consume a few computational overhead [wu2023parameter], but the incurred latency is non-negligible. As revealed in recent progresses [luo2023cheap, gao2023llama], the additional branches added by adapters inevitably slow down inference although they only brings a few of FLOPs [luo2023cheap]. The same as shown in Fig.1, the latency caused by adapters takes up to 22.3% inference time of Adapter [sung2022vl]. Conclusively, correctly judging the redundancy of MLLMs to achieve true speed-up is one main obstacle of PCETL.
@@
To handle these challenges, we propose a novel approach for the parameter and computation efficient tuning of MLLMs in this paper, termed Effective Attention Skipping (EAS). In particular, we first investigate the redundancy of different components of MLLMs, and show that skipping MHA can reduce more computation overhead without impeding performance, as shown in Fig. 1-(b). The principle of EAS is to only skip the redundant MHAs for better efficiency and performance. To avoid the latency caused by adapters [wu2023parameter], we also equip EAS with an innovative parameter efficient module called propagation-of-information adapter (PIA), which can not only replace the skipped MHAs for adaption, but also can be seamlessly re-parameterized into FFNs for zero extra cost during inference. With these innovative designs, EAS can help MLLMs better approach the targets of PCETL on vision-language tasks, i.e., true model acceleration with a small number of trainable parameters.
@@
To validate EAS, we first apply it to two MLLMs, i.e., LLaVA [llava-1.5] and LaVIN [luo2023cheap], and conduct extensive experiments on two out-domain benchmarks, Slake and AID, and the open-set multi-modal question answer benchmark, namely ScienceQA [lu2022learn]. To align DAS [wu2023parameter], we also generalize EAS to a representative VL pre-trained model called METER [conf/cvpr/DouXGWWWZZYP0022] on three VL benchmarks, i.e., VQA2.0 [conf/cvpr/GoyalKSBP17], NLVR2 [conf/acl/SuhrZZZBA19] and Flickr30K [journals/ijcv/PlummerWCCHL17]. The experimental results show that EAS can not only retain high performance and parameter efficiency against existing PETL and PCETL methods [zhang2023llama, luo2023cheap, wu2023parameter], but also significantly speed up inference speed on downstream tasks. For instance, EAS can improve the inference speed by $1.20\times$ to the default LLaVA [liu2023visual] without performance degradations.
Overall, our contributions are three-fold:
@@
[…]
Figure 3: Illustrations of the proposed Effective Attention Skipping (EAS). (a) The architecture of propagation-of-information adapter (PIA). PIA uses a multi-path design for up- and down-samplings, which can act a role of MHA in information exchange for MLLMs. (b) The deployment of PIA. PIA can serve to replace the skipped MHA as a parameter efficient method for task adaption. After training, its parameters can be re-parameterized into FFN, incurring no extra latency. (c) The process of attention redundancy evaluation. EAS also adopts a $k$-armed bandit based algorithm for the automatic redundancy evaluation on MHAs of MLLMs. After evaluation, we skip the redundant MHAs with PIAs.
@@ 2.1 Vision-language pre-trained models
Recent years have witnessed the rapid development of vision-language pre-trained (VLP) models [li2019visualbert, lu2019vilbert, su2019vl, cho2021unifying, conf/icml/WangJLZCYL22] on various vision-language tasks [journals/ijcv/GoyalKASBP19, lu2022learn, journals/ijcv/JiangLZ24]. To facilitate vision-language alignment and learn generalized multi-modal representations, VLP models also adopt self-supervised objects and conduct pre-training on massive image-text pairs, such as masked language modeling [devlin2018bert, liu2019roberta], masked image modeling [conf/iclr/Bao0PW22, conf/icml/ChenRC0JLS20, conf/iclr/DosovitskiyB0WZ21] and image-text matching [conf/cvpr/DouXGWWWZZYP0022, conf/icml/KimSK21].
With the advent of large language models (LLMs) [zhang2205opt, touvron2023llama, openai2023gpt], recent advance [luo2023cheap, liu2023visual, li2023blip, dong2021dual] focus on extending these LLMs to vision-language tasks [guo2024benchmarking, marino2019ok, conf/cvpr/GoyalKSBP17], i.e., multi-modal large language models (MLLMs). For instance, BLIP-2 [conf/icml/0008LSH23] introduces a QFormer, to bridge the gap between vision and language modalities, selecting relevant visual semantics as additional tokens for LLMs. MINI-GPT4 [journals/corr/abs-2304-10592] uses a projection layer to map visual features into the language model. LLaVA [conf/nips/LiuLWL23a] also follows this principle, but proposes a carefully designed training scheme in addition. Meanwhile, some efforts are also devoted to parameter-efficiently transfer learning of LLMs to vision-language tasks [goyal2017making, lu2022learn]. These methods insert lightweight modules like adapter [zhang2023llama] for downstream adaption instead of fully tuning LLMs [hu2022lora]. Meanwhile, the use of visual tokens greatly exacerbates the already high computation of MLLMs [liu2023visual, zhang2023llama]. In this paper, we focus on both parameter and computation efficient tuning for MLLMs.
@@ 2.2 Parameter and Computation Efficient Tuning
Parameter-efficient transfer learning (PETL) is proposed to save the training and storage costs of LLMs [brown2020language, li2021prefix, hu2022lora]. The principle of PETL is to transfer the pre-trained models to downstream tasks with only a small number of trainable parameters [brown2020language, li2021prefix], and its methodologies can be roughly divided into three main categories, i.e. prompt tuning [jia2022visual, zhou2022learning, journals/ijcv/BulatT24], adapter [karimi2021compacter, gao2023clip] and the re-parameterized methods [hu2022lora, luo2023towardsefficient]. These methods mainly focus on efficiently transferring models to the downstream tasks. PEQA [kim2023memory] proposes a simple method that only updates quantization scales with quantized LLMs to reduce the requirements of memory during the training. QLoRA [dettmers2023qlora] quantizes and freezes the original model, and uses LoRA to transfer the model to downstream tasks. HydraLoRA [tian2024hydralora] uses an asymmetric structure based on LoRA to reduce the need for domain expertise in transferring. Recently, [wu2023parameter] propose a new learning task called parameter and computation efficient transfer learning (PCETL), which further requires to reduce model redundancy based on the setting of PETL. [wu2023parameter] also propose a strong baseline called Dynamic Architecture Skipping (DAS) for PCETL, which evaluates the redundancy of Transformer layers by a reinforcement learning method and replaces the most redundant layers with adapters [sung2022vl]. Motivated by [hu2022lora, wu2023parameter], we also focus on PCETL for MLLMs in this paper. We propose a granular exploration for MLLMs, and address the latency caused by using adapters for skip connections.
2.3 Improving Efficiency in MLLMs
The inference efficiency of the MLLM has also attracted much attention from researchers. In recent years, many works focus on improving the inference efficiency of MLLM, with the main methods including simplifying the structure of models [tong2024flashsloth, zhou2024tinyllava, zhu2024llava] and reducing the length of the inference sequence [zhang2024cls, han2024rethinking]. In the way to reduce the inference sequence, FastV [chen2024image] uses a metric to evaluate the importance of visual tokens and prunes half of them to reduce computation overhead. FitPrune [ye2025fit] removes visual tokens without affecting performance by fitting changes in the attention matrix. However, these methods mainly focus on reducing the number of visual tokens, and efficiency improvement is mainly in the prefilling stage. In terms of structural design, recent MLLMs like MoE-LLaVA [moe_llava] selectively activate expert parameters through the MoE structure. Mono-InternVL [luo2024mono] splits the parameters into text and vision experts to improve the efficiency. And RoE [wu2024routing] sparses the model through selectively activating the layers in inference. However, these methods rely on fully tuning MLLMs. In this paper, we achieve both parameter and computation efficiency in transferring MLLMs to the downstream task.
@@ 3 Preliminary
We first recap the principle of parameter and computation efficient transfer learning (PCETL) [wu2023parameter] for multi-modal large language models (MLLMs). Concretely, given an MLLM $G(\cdot|\theta)$, which consists $n$ Transformer layers, and its whole parameters denoted as $\mathbf{\theta}=\{\theta_{1},\theta_{2},...,\theta_{n}\}$, the objective of PCETL can be defined by
@@
From Eq.1, we can see that PCETL needs to reduce parameter cost during adaption, just like PETL [li2021prefix, hu2022lora, sung2022vl]. Meanwhile, it also requires practitioners to speed up inference by evaluating and removing the redundant modules. Compared with previous model compression and acceleration tasks [krishnamoorthi2018quantizing, meng2020pruning, dong2021attention], one difference of PCETL is that the original model weights should keep unchanged, so as to facilitate more task adaptions, which is essential for the giant models that are expensively pre-trained with massive data.
[wu2023parameter] also propose a strong baseline called dynamic architecture skipping (DAS), which considers the entire Transformer layer for redundancy evaluation. As discussed above, the components in MLLMs often serve different roles, i.e., MHA and FFN. In this case, the granular evaluations are beneficial for PCETL. Besides, the skip connections in DAS are aided by adapters [houlsby2019parameter, sung2022vl] which also incur non-negligible latency during the inference [wu2023parameter, hu2022lora].
@@ 4.1 Overview
Concretely, given an MLLMs $G(\cdot|\theta)$, EAS aims to remove its redundant components for downstream task adaption. Different to DAS [wu2023parameter], the evaluation target of EAS is more specific, i.e., MHAs of a MLLM. Thus, the objective can be defined by
@@
where $\bm{\theta}^{A}_{\mathcal{K}}$ represent a subset of MHA modules and $\bm{\theta}^{F}$ denote all FFN modules. In practice, similar to DAS [wu2023parameter], we first conduct a reinforcement learning based redundancy evaluation on MHAs for downstream tasks. Afterwards, given the redundancy scores, we determine which MHAs to skip, and replace them with an adapter-based connection.
However, adapters will account for excessive inference time, as shown in Fig. 1-(a). The widely used re-parameterized methods like LoRA [hu2022lora] can approximate the QKV projections but cannot replace the entire MHA for skip connections [luo2023towardslanguage].
@@ 4.2 Propagation-of-Information Adapter
When redundant MHAs are identified, the next key step is to effectively skip them without additional latency. To approach this target, we propose a novel re-paremeterizable adapter, termed Propagation-of-Information Adapter (PIA).
As shown in Fig. 3-(a), PIA also adopts a bottleneck structure to scale down the hidden dimension of the inputs, akin to existing adapters [sung2022vl, luo2023towardsefficient], thereby achieving a low-rank approximation of full tuning [hu2022lora, wu2023approximated]. However, it also differs in its inner path designs.
@@
where $f_{d1}(\cdot)$ and $f_{d2}(\cdot)$ are the two linear projections, and $avg(\cdot)$ refers to average pooling. $\mathbf{H}^{(i)}\in\mathbb{R}^{n\times r}$ has a much smaller dimension than $\mathbf{X}^{(i)}$, i.e., $r\ll d$. Besides, with the information propagation through the average pooling, the skipping of MHA can lead to less impact.
@@
During up-sampling, we also use two separate linear projections for the hidden features, and adopt a path router for their weighted combination. Thus, the output $\mathbf{X}^{(i)^{\prime}}$ is obtained by
@@
$\mathrm{where}$ $\ \ [\alpha_{1},\alpha_{2}]=\mathrm{router}(avg(\mathbf{X}^{(i)})).$
Here, $f_{u1}(\cdot)$ and $f_{u2}(\cdot)$ are two linear projections. Their weights will be given by the router in the prefilling stage. And $\mathrm{router}(\cdot)$ is the routing function defined by
@@
Overall, the multi-path design can still hold parameter efficiency via low-dimensional projections, acting the role of MHA in information exchange [vaswani2017attention].
Re-parameterization. With the effective structure, PIA also needs to re-parameterize its weights into the models to further improve computation efficiency. Here, we embed PIA into FFN, and avoid the skip connections [wu2023parameter] to retain network complexity.
@@
To re-parameterize PIA into the nearby projection weights of FFN [hu2022lora, luo2022towards], we still need to convert PIA into one linear layer:
@@
Here, $\mathbf{W}_{p}$ is the approximated low-rank weight matrix, and it then can be re-parameterized into the nearby weight matrices of FFN, like [hu2022lora, luo2023towardsefficient].
Note that, the average feature in Eq. 4, i.e. $\hat{\mathbf{x}}^{(i)}=avg(\mathbf{X}^{(i)})$, and the path router are example-dependent. In this case, the re-parameterization of PIA will be executed after the first decoding step of MLLMs. Considering the fact that most MLLMs often needs to decode a long sequence [luo2023cheap, zhang2023llama], PIA can still save massive computation during inference.